# Semantic Parallelism: Redefining Efficient MoE Inference via Model-Data Co-Scheduling

### v4

@@ 1 Introduction
To address the memory demands of large-scale MoE deployment and leverage aggregate memory bandwidth, modern inference engines such as SGLang Zheng et al. (2024) and vLLM Kwon et al. (2023) employ expert parallelism (EP), whereby experts are distributed across devices. Attention layers are typically parallelized via data parallelism (DP) or tensor parallelism (TP). While EP enables parallel computation of experts across GPUs, it introduces significant communication overhead: intermediate activations must be dispatched from the gating module on a source GPU to the destination GPUs hosting the routed experts, and later combined back after expert computation. These operations often result in cluster-wide any-to-any token shuffling, typically implemented via two all2all collective operations (e.g., NCCL’s all2all).
Figure 1: Latency breakdown for DeepSeek-V2-Lite inference over a single MoE layer. Hardware: 8-GPU (96GB) server with fast inter-GPU network (900GB/s).
@@ 2 Background
MoE Load-balancing and Experts Re-grouping. Lina Li et al. (2023) probes the variation of expert hotness and allots non-uniform expert replicas to achieve load-balanced expert computation. Similar studies Huang et al. (2023) exist to pursue expert load balancing and mitigate other sources of MoE computing inefficiencies. EPS-MoE Qian et al. (2025) optimizes the computation of MoE FeedForward Network (FFN) modules by dynamically selecting the best backend implementation of GroupGemm and DenseGemm. DeepSeek also adopts EPLB (expert-parallelism load balancing) in its real-world deployment DeepSeek-AI (2024b). ExFlow Yao et al. (2024) exploits the affinity between experts across adjacent layers to reduce remote routing and collocate closely related experts. Exflow only considers the model scheduling for MoE models, and requires a heavy allgather before the execution of each model layer, which significantly incurs memory pressure and extra communication overhead. MoETuner Go and Mahajan (2025) optimizes the MoE model serving by finding an optimal expert placement strategy to minimize inter-device communication.
@@ 3.2 Predicting Expert Routing Path
Context-Independent Token Activation Prediction. We observe that despite the theoretical dependence of expert routing on contextual semantics (as expressed by the gating function $G_{L}(h_{L,j})$), in practice, tokens exhibit strong context-independent affinities to specific experts. This enables effective prediction based solely on token identity. Through offline profiling on datasets such as Sharegpt using models including DeepSeek-V2-lite and Qwen3-30B-A3B (see Appendix Figure 6), we construct a token-to-expert activation table $\mathbf{T}^{(L)}\in\mathbb{N}^{t\times N^{(L)}}$ for each MoE layer $L$, where $\mathbf{T}^{(L)}_{j,k}$ counts how frequently token $x_{j}$ activates expert $E_{k}^{(L)}$. The corresponding routing probability is: $\Pr(E^{(L)}_{k}|x_{j})=\mathbf{T}^{(L)}_{j,k}/\sum_{k=1}^{N^{(L)}}\mathbf{T}^{(L)}_{j,k}$. For efficient online inference, these probabilities are tabulated in a token-to-expert confidence table $\bm{\mathcal{C}}_{p}\in\mathbb{R}^{t\times N}$. Out-of-vocabulary tokens are handled via nearest-neighbor matching in the embedding space.
@@ 4.1 Experimental Environments
We evaluate Sem-MoE on an 8-GPU server, representing commercial GPU servers unified for both training and inference, which are configured with 96GB-HBM per GPU and fast homogeneous interconnects. GPUs inside a server can communicate with each other at a premium bandwidth (900GBps). The server is equiped with two 44-core Intel CPUs and 2TB DDR5 memory.
@@ 4.3 Baselines and Performance Metrics
MoETuner: MoETuner Go and Mahajan (2025) is an optimization framework that enhances MoE model serving performance by finding an optimal expert placement strategy. It addresses critical bottlenecks in expert parallelism, namely imbalanced token processing loads across GPUs and skewed inter-GPU communication, which lead to significant tail latency. The core of MoETuner is an Integer Linear Programming (ILP) formulation that leverages predictable token routing dependencies across layers. It jointly optimizes expert-to-GPU assignments to balance computational workloads and minimize communication costs, thereby reducing end-to-end execution time. We embed MoETuner into SGLang as a comparable baseline.
@@
Table 1: Attention-TP Scenario: TTFT and E2E latency under different
@@ 4.4 End-to-End Inference Performance
Attention-TP Scenario. For the attention-TP scenario, there is no scheduler for a single inference instance, as different TP ranks receive the same input. Meanwhile, latency becomes a main concern in TP settings. Therefore, we set the request rate to 1 req/s and vary the input sequence length to observe the TTFT and end-to-end latency directly as shown in Table 1. For Deepseek-V2-Lite, Sem-MoE outperforms the best baseline in TTFT by 12.21%, 10.60%, and 18.89% under input lengths of 256, 512, and 1024. For Qwen3-30B-A3B, the performance optimization ratio is 17.16%, 24.90%, 3.80%. Thanks to the load-balance and inter-layer communicaiton optimizing effect, MoETuner gains performance improvement in some cases, yet may slow down in the other cases. Model-data collaborative scheduling can bring holistic performance boosting in all tested scenarios. The speedup of TTFT also translates to the shrinking of end-to-end inference latency, just as shown in Table 1. We would delve into the execution of MoE layers to analyze the rationale behind the breakdown of the inference speedup.
@@ 5 Conclusion
The communication overhead of expert parallelism renders a significant bottleneck in serving large-scale MoE models. We present Sem-MoE, which can proactively and losslessly trim EP’s all2all communication volume via model-data collaborative scheduling, leveraging the intrinsic affinity between model experts and input tokens. Experiments show that Sem-MoE can significantly reduce communication overhead and boost inference throughput under differently specified SLO constraints, both in attention-DP and attention-TP scenarios.
@@ A.1 Impact of Request Semantics on Expert Activation
A.2 Intra-layer bi-clustered token-expert Conjugacy
@@
The above examples illustrate the tabularized relationship between tokens and experts. We argure that, simple conditional probability model can work to predict the activation path of tokens by combining the above intra-layer conjugacy and/inter-layer affinity. We first calculate the kurtosis [^3] Kurtosis is a measure of the tailedness of a distribution. High Kurtosis indicates a token favors several fixed experts during multiple occurrences. of each token’s activation map. The right picture of Figure 6(b) shows most of the kurtosis values are higher than 8, indicating that the to-route experts for each token concentrate on a narrowed set, regardless of the context. We further use partial (25%) of the profile dataset to calculate each tokens’ most routed top-k experts. Then the static top-k experts are used to predict the tokens activation using the left part (75%) of the profile dataset, achieving a 96.3% precision and a 78.8% F1 score. The right part of Figure 6(c) predicts the next-layer-activation via looking back the prior layers. As we know more about the previous activation sequence at layer $L$, we can predict the activation at layer $L+1$ with higher confidence (about 70% when looking back 5 layers).
@@ 53       $maks[cls_{r}]$ $\leftarrow$ 0
Leveraging the intrinsic conjugacy between tokens and experts in MoE models to achieve token-expert co-dispatching, may help reduce communication volume and boost distributed parallel inference. First, similar experts can be pre-grouped together at deployment time based on pre-profiled and modeled affinity. Second, on the fly, individual tokens can be re-shuffled and re-batched to the GPU devices that host the expert groups whose member experts have largest modeled conjugacy with the token. Such co-scheduling aims to avoid scattered, cross-device token-expert activations, or equivalently, maximize the probability of intro-device, local activations. Figure 6(d) shows an micro-benchmark experiment, testing the all2all latency under different local activation rate using the nccl all2allv API. With the local activation rate ($\alpha$) varying from 0.2 to 0.9, the latency decreases gradually, showing the performance gains with improved expert-token co-scheduling.

### v5

# arXiv:2503.04398v5[cs.LG] 27 Feb 2026

@@ 1 Introduction
To address the memory demands of large-scale MoE deployment and leverage aggregate memory bandwidth, modern inference engines such as SGLang Zheng et al. (2024) and vLLM Kwon et al. (2023) employ expert parallelism (EP), whereby experts are distributed across devices. Attention layers are typically parallelized via data parallelism (DP) or tensor parallelism (TP). While EP enables parallel computation of experts across GPUs, it introduces significant communication overhead: intermediate activations must be dispatched from the gating module on a source GPU to the destination GPUs hosting the routed experts, and later combined back after expert computation. These operations often result in cluster-wide any-to-any token shuffling, typically implemented via two all2all collective operations (e.g., NCCL/HCCL’s all2all).
Figure 1: Latency breakdown for DeepSeek-V2-Lite inference over a single MoE layer. Hardware: 8-GPU server with fast inter-GPU network (specialized connection with over 400GB/s bandwidth).
@@ 2 Background
MoE Load-balancing and Experts Re-grouping. Lina Li et al. (2023) probes the variation of expert hotness and allots non-uniform expert replicas to achieve load-balanced expert computation. Similar studies Huang et al. (2023) exist to pursue expert load balancing and mitigate other sources of MoE computing inefficiencies. EPS-MoE Qian et al. (2025) optimizes the computation of MoE FeedForward Network (FFN) modules by dynamically selecting the best backend implementation of GroupGemm and DenseGemm. DeepSeek also adopts EPLB (expert-parallelism load balancing) in its real-world deployment DeepSeek-AI (2024b). ExFlow Yao et al. (2024) exploits the affinity between experts across adjacent layers to reduce remote routing and collocate closely related experts. Exflow only considers the model scheduling for MoE models, and requires a heavy allgather before the execution of each model layer, which significantly incurs memory pressure and extra communication overhead. MoETuner Go & Mahajan (2025) optimizes the MoE model serving by finding an optimal expert placement strategy to minimize inter-device communication.
@@ 3.2 Predicting Expert Routing Path
Context-Independent Token Activation Prediction. We observe that despite the theoretical dependence of expert routing on contextual semantics (as expressed by the gating function $G_{L}(h_{L,j})$), in practice, tokens exhibit strong context-independent affinities to specific experts. Table 1 presents three metrics illustrating the activation concentration within DeepSeek-V2-Lite. By profiling the model on the ShareGPT Datasets (2023) dataset, we record the intermediate top-k activation statistics. We then calculate the F1-score for predicting the activated experts in each layer by relying solely on the most frequently activated (hottest) top-k experts. Additionally, we compute the cumulative hotness of these top-k experts alongside the maximum hotness observed among the remaining non-top-k experts. For all evaluated metrics, the P25, P50, and P75 quantiles are reported.
For any specific token, the distribution of expert activation across diverse contexts is remarkably skewed yet stable: a given token consistently routes to a narrow, static subset of top-k experts, leaving the majority of experts largely dormant. Throughout all layers of DeepSeek-V2-Lite, the median (p50) cumulative hotness of the top-k experts ranges from 0.833 to 0.976. In contrast, the maximum hotness recorded among the cold experts remains negligible—hovering near 0.05 at p50—which highlights a robust and pronounced disparity in activation. Finally, § A.2 in the Appendix expands upon this phenomenon across more scenarios, confirming its broad generalizability.
The above phenomenon enables effective prediction based solely on token identity. Through offline profiling on datasets such as Sharegpt using models including DeepSeek-V2-lite and Qwen3-30B-A3B, we construct a token-to-expert activation table $\mathbf{T}^{(L)}\in\mathbb{N}^{t\times N^{(L)}}$ for each MoE layer $L$, where $\mathbf{T}^{(L)}_{j,k}$ counts how frequently token $x_{j}$ activates expert $E_{k}^{(L)}$. The corresponding routing probability is: $\Pr(E^{(L)}_{k}|x_{j})=\mathbf{T}^{(L)}_{j,k}/\sum_{k=1}^{N^{(L)}}\mathbf{T}^{(L)}_{j,k}$. For efficient online inference, these probabilities are tabulated in a token-to-expert confidence table $\bm{\mathcal{C}}_{p}\in\mathbb{R}^{t\times N}$. Out-of-vocabulary tokens are handled via nearest-neighbor matching in the embedding space.
@@
Table 1: Empirical study of token activate patterns based on DeepSeek-V2-Lite.
Layer ID F1-score Cum. Hotness (Top-k) Max. Hotness (Non-top-k)
p25 p50 p75 p25 p50 p75 p25 p50 p75
1 1.000 1.000 1.000 0.833 1.000 1.000 0.000 0.042 0.071
6 0.833 1.000 1.000 0.767 0.917 1.000 0.000 0.056 0.083
11 0.833 1.000 1.000 0.722 0.889 1.000 0.000 0.056 0.083
16 0.667 0.833 1.000 0.708 0.889 1.000 0.000 0.053 0.083
21 0.667 0.833 1.000 0.667 0.833 1.000 0.000 0.050 0.074
26 0.667 0.833 1.000 0.667 0.833 1.000 0.000 0.051 0.083
@@ 4.1 Experimental Environments
We evaluate Sem-MoE on an 8-GPU server, representing commercial GPU servers unified for both training and inference, which are configured with 96GB-HBM per GPU and fast homogeneous interconnects. GPUs inside a server can communicate with each other at a premium bandwidth (specialized network with over 400GB/s bandwidth). The server is equiped with two 44-core Intel CPUs and 2TB DDR5 memory.
@@ 4.3 Baselines and Performance Metrics
MoETuner: MoETuner Go & Mahajan (2025) is an optimization framework that enhances MoE model serving performance by finding an optimal expert placement strategy. It addresses critical bottlenecks in expert parallelism, namely imbalanced token processing loads across GPUs and skewed inter-GPU communication, which lead to significant tail latency. The core of MoETuner is an Integer Linear Programming (ILP) formulation that leverages predictable token routing dependencies across layers. It jointly optimizes expert-to-GPU assignments to balance computational workloads and minimize communication costs, thereby reducing end-to-end execution time. We embed MoETuner into SGLang as a comparable baseline.
@@
Table 2: Attention-TP Scenario: TTFT and E2E latency under different input lengths.
@@ 4.4 End-to-End Inference Performance
Attention-TP Scenario. For the attention-TP scenario, there is no scheduler for a single inference instance, as different TP ranks receive the same input. Meanwhile, latency becomes a main concern in TP settings. Therefore, we set the request rate to 1 req/s and vary the input sequence length to observe the TTFT and end-to-end latency directly as shown in Table 2. For Deepseek-V2-Lite, Sem-MoE outperforms the best baseline in TTFT by 12.21%, 10.60%, and 18.89% under input lengths of 256, 512, and 1024. For Qwen3-30B-A3B, the performance optimization ratio is 17.16%, 24.90%, 3.80%. Thanks to the load-balance and inter-layer communication optimizing effect, MoETuner gains performance improvement in some cases, yet may slow down in the other cases. Model-data collaborative scheduling can bring holistic performance boosting in all tested scenarios. The speedup of TTFT also translates to the shrinking of end-to-end inference latency, just as shown in Table 2. We would delve into the execution of MoE layers to analyze the rationale behind the breakdown of the inference speedup.
Extensive Experiments. We also conduct performance evaluation using one additional model (Moonlight-16B Liu et al. (2025)), and the results are consistent with the above observations, which are provided in § C.2.
@@ 4.6 Algorithm Evaluation
We also evaluate the cross-dataset zero-shot transfer performance of the predictor model, proving the robustness of the model-data affinity and the generalizability of the scheduling policy, which can be referred to § C.1.
@@ 5 Conclusion
The communication overhead of expert parallelism renders a significant bottleneck in serving large-scale MoE models. We present semantic parallelism and Sem-MoE, which can proactively and losslessly trim EP’s all2all communication volume via model-data collaborative scheduling, leveraging the intrinsic affinity between model experts and input tokens. Experiments show that Sem-MoE can significantly reduce communication overhead and boost inference throughput under differently specified SLO constraints, both in attention-DP and attention-TP scenarios.
@@ A.1 Impact of Request Semantics on Expert Activation
A.2 Token-expert Conjugacy
Intra-layer token-expert affinity
@@ A.2 Token-expert Conjugacy
The above examples illustrate the tabularized relationship between tokens and experts. We argue that, simple conditional probability model can work to predict the activation path of tokens by combining the above intra-layer conjugacy and/inter-layer affinity. We first calculate the kurtosis [^3] Kurtosis is a measure of the tailedness of a distribution. High Kurtosis indicates a token favors several fixed experts during multiple occurrences. of each token’s activation map. The right picture of Figure 6(b) shows most of the kurtosis values are higher than 8, indicating that the to-route experts for each token concentrate on a narrowed set, regardless of the context. We further use partial (25%) of the profile dataset to calculate each tokens’ most routed top-k experts. Then the static top-k experts are used to predict the tokens activation using the left part (75%) of the profile dataset, achieving a 96.3% precision and a 78.8% F1 score. The right part of Figure 6(c) predicts the next-layer-activation via looking back the prior layers. As we know more about the previous activation sequence at layer $L$, we can predict the activation at layer $L+1$ with higher confidence (about 70% when looking back 5 layers).
@@ 53       $maks[cls_{r}]$ $\leftarrow$ 0
Leveraging the intrinsic conjugacy between tokens and experts in MoE models to achieve token-expert co-dispatching, may help reduce communication volume and boost distributed parallel inference. First, similar experts can be pre-grouped together at deployment time based on pre-profiled and modeled affinity. Second, on the fly, individual tokens can be re-shuffled and re-batched to the GPU devices that host the expert groups whose member experts have largest modeled conjugacy with the token. Such co-scheduling aims to avoid scattered, cross-device token-expert activations, or equivalently, maximize the probability of intra-device, local activations. Figure 6(d) shows an micro-benchmark experiment, testing the all2all latency under different local activation rate using the nccl all2allv API. With the local activation rate ($\alpha$) varying from 0.2 to 0.9, the latency decreases gradually, showing the performance gains with improved expert-token co-scheduling.
@@ 18 $\mathcal{B}$ $\leftarrow$ $\mathcal{B}\left[r\_shf\_indices\right]$
Appendix C Extensive Experiments
C.1 Cross-dataset Validation of Expert Routing Prediction
To assess the generalization of our method, we conducted cross-domain experiments using two large language models: DeepSeek-V2-Lite and Qwen3-30B. Specifically, we trained Sem-MoE’s activation predictor on a single source dataset and evaluated the quality of its token-expert co-scheduling decisions, measured by the Local Activation Rate (LAR), on the other two unseen, out-of-distribution target datasets. Note that LAR measures the proportion of tokens routed to local experts, reflecting a reduced communication volume to experts on remote devices. Therefore, a higher LAR indicates lower cross-device communication overhead—and is better.
The results reveal robust zero-shot transfer performance for our method.
Table 3: Cross-dataset evaluation of zero-shot transfer for DeepSeek-V2-Lite
Method LAR (p50) ShareGPT Lmsys-Chat-1M MMLU
Sem-MoE trained on ShareGPT 46.49% 41.25% 41.55%
Sem-MoE trained on Lmsys-Chat-1m 40.68% 47.19% 41.40%
Baseline: SGLang 24.98% 25.01% 24.97%
Table 4: Cross-dataset evaluation of zero-shot transfer for Qwen3-30B
Method LAR (p50) ShareGPT Lmsys-Chat-1M MMLU
Sem-MoE trained on ShareGPT 46.88% 39.28% 35.30%
Sem-MoE trained on Lmsys-Chat-1m 36.57% 43.61% 33.37%
Baseline: SGLang 25.00% 25.01% 25.00%
As shown in Table 3, for DeepSeek-V2-Lite, the predictor trained on ShareGPT achieves an in-distribution LAR of 46.49%. When applied without retraining or fine-tuning to Lmsys-Chat-1M, it still attains 41.25% LAR—only a modest drop from the best in-domain result on that dataset (47.19%, achieved when training directly on Lmsys-Chat-1M). More importantly, this out-of-distribution performance is  1.65× higher than the SGLang’s default scheduling setting (25.01%). Similarly, on MMLU—a domain with very different content and structure—the same ShareGPT-trained predictor yields 41.55% LAR, far surpassing SGLANG defaults (24.97%) and remaining close to in-distribution levels.
As shown in Table 4, the trend is consistent for Qwen3-30B: the ShareGPT-trained predictor achieves 39.28% LAR on Lmsys-Chat-1M and 35.30% on MMLU, compared to peak in-domain scores of 43.61% and 36.57% (when trained on Lmsys), and both are substantially above the SGLang baseline ( 25%). A similar trend holds when transferring reservedly from Lmsys-Chat-1M to ShareGPT and MMLU.
These results demonstrate that when fed with real-world, representative datasets, Sem-MoE tends to capture the invariability in token-routing patterns that remain effective across different linguistic styles, task types, and knowledge domains. While training on the target distribution offers marginal improvements, it is not required except for extreme performance pursuits.
C.2 Throughput Improvement of Moonlight-16B
To address the issue of model diversity, we evaluate the Moonlight-16B model. The results align closely with the performance trends observed for DeepSeek-V2-Lite and Qwen3-A30B, as reported in the § 4. In summary, our approach Sem-MoE, consistently outperforms SGLANG (MoETuner): it achieves an average performance improvement of 1.12× (1.22×) for prefill and 1.10× (1.14×) for end-to-end inference, respectively. Detailed results are provided in Table 5 and Table 6.
Table 5: Throughput Optimization Effect under TTFT SLO
Dataset Throughput (Req/s) SGLang MoETuner Sem-MoE
Speedup
v.s. SGLang
Speedup
v.s. MoETuner
Lmsys-Chat-1M 32.7 28.6 38.6 1.18x 1.35x
MMLU 32.2 30.1 38.6 1.20x 1.28x
ShareGPT 2.9 2.8 2.9 1.00x 1.02x
Average 22.6 20.5 26.7 1.12x 1.22x
Table 6: Throughput Optimization Effect under E2E latency SLO
Dataset Throughput (Req/s) SGLang MoETuner Sem-MoE
Speedup
v.s. SGLang
Speedup
v.s. MoETuner
Lmsys-Chat-1M 37.2 35.5 38.6 1.04x 1.09x
MMLU 37.9 35.3 38.6 1.02x 1.10x
ShareGPT 1.9 1.9 2.4 1.23x 1.23x
Average 25.7 24.2 26.5 1.10x 1.14x